# An Ecosystem for Ontology Interoperability

### v2

@@ Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
@@ I Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a logic-based expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain. There is no guarantee that ontologies will agree on the same conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies. For example, it is common for ontologies to encompass a range of different concepts. A person can be described using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee). Moreover, the terminology used in ontologies can be different. It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
@@ II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2]. Fig. 1 illustrates that ODPs use the star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
@@
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies. They are static and are not prone to capturing changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
@@ II-B Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. Fig. 2 shows that OM&OV conform to the pairwise matching schema. This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
@@
OM&OV are typically resource-consuming. The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
@@ II-C Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) [5]. Fig. 3 shows that the OCKGs comply with the star-based matching schema. While ontologies comply with KG(s), their reference mappings can be derived from the concepts pointing to the same data instance.
@@
Building OCKGs is not easy. This process requires extensive workloads to find the alignments between different ontologies, so that each KG can be modelled with different ontologies to determine which ontology fits the KG(s) better.
@@ III An Ecosystem
Fig. 4 illustrates an overview of the ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
@@ IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is recommended to build an ontology based on ODPs. Several ODPs can be used in this case, as illustrated in Fig. 5.
@@
While all of these ODPs can effectively capture the main concepts in the example, they have differences that need to be considered.
@@
    The definition of the concept can be different. For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
@@ IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN. As illustrated in Fig. 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. If we consider SOSA to be an updated version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [12]. OM&OV tasks can be classified into two main categories.
@@ IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it remains challenging to determine which ontology to use in the real world. Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
@@
To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings (retrieved from [13, 14, 15]).
@@ V Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Introduced in 2023, ontology-compliant KGs aim to explore the interoperability between ontology and its complementary KGs.
@@
The source code, data, and/or other artifacts used in this study have been made available at https://github.com/qzc438-research/ontology-interoperability.
@@ Acknowledgment
The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v3

@@ Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability and data integration in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
@@ I Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a class schema expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a person, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application. There is no guarantee that ontologies in the same domain agree on the conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies. For example, it is common for ontologies to encompass a range of different concepts, such as describing a person using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee). Moreover, the terminology used in ontologies can be different. It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
In this paper we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
@@ II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2]. Fig. 1 illustrates that ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
@@
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly hand-crafted on the basis of existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
@@ II-B Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. Fig. 2 shows that OM&OV conform to a pairwise matching schema. This means that each mapping is a pairwise process to find the correspondences between two ontologies.
@@
OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in each of the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
@@ II-C Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Fig. 3 shows that OCKGs comply with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
@@
Building OCKGs is not easy. To determine which candidate ontology fits the KG(s) better, the KG needs to be modelled with each candidate. This process itself is essentially building an alignment of the candidate ontologies, informed by the KG under study.
@@ III An Ecosystem
Fig. 4 illustrates an overview of an ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve interoperability.
@@ IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Fig. 5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives. While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown in the following.
@@
On the other hand, differences in ontology characteristics, such as those listed below, need to be addressed in practice.
@@
    The definition of the concepts can be different. For example, temperature is defined as a property of the room in both Observation Pattern and Measurement Pattern, but in the Event Pattern, it is considered a quality measure of the room. Activity Pattern defines the temperature as a member of the room.
@@ IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Fig. 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
@@ IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts within the given KG.
@@
To model these two new concepts and adequately integrate them with the other parts of a KG, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings (retrieved from [13, 15, 16]).
@@ V Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Ontology-compliant knowledge graphs (OCKGs) aim to explore the interoperability between an ontology and its complementary KGs.
@@
The source code, data, and/or other artifacts used in this study are available at https://github.com/qzc438-research/ontology-interoperability.
@@ Acknowledgment
The author is supervised by Kerry Taylor (Australian National University), Weiqing Wang (Monash University), Jing Jiang (Australian National University), Subbu Sethuvenkatraman (Commonwealth Scientific and Industrial Research Organisation), Qing Wang (Australian National University), and Alex Potanin (Australian National University). The author thanks the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.